# Mass Modeling and Kinematics of Galaxy Clusters in Modified Gravity

### v1

@@ Abstract
The chameleon screening mechanism has been constrained many a time using dynamic and kinematic galaxy cluster observables. Current constraints are, however, insensitive to different mass components within galaxy clusters and have been mainly focused on a single mass density profile, the Navarro-Frenk-While mass density model. In this work, we extend the study of the Chameleon screening mechanism in galaxy clusters by considering a series of mass density models, namely: generalized-Navarro-Frenk-While, b-Navarro-Frenk-While, Burket, Isothermal and Einasto. The coupling strength ($\beta$) and asymptotic value of the chameleon field ($\phi_{\infty}$) are constrained by using kinematics analyses of simulated galaxy clusters, generated both assuming General Relativity and a strong chameleon scenario. By implementing a Bayesian analysis we comprehensively show that the biases introduced due to an incorrect assumption of the mass model are minimal. Similarly, we also demonstrate that a spurious detection of evidence for modifications to gravity is highly unlikely when utilizing the kinematics of galaxy clusters.
@@ 1 Introduction
Among the plethora of possibilities, Chameleon field theories [35, 36, 37] deliver some of the most promising results. In the latter, the environment-dependent effects of gravity are achieved by a non-minimally coupling between an additional scalar field (aka chameleon field) and matter, with the screening achieved by the scalar fields’ potential dependency on the local energy density[38]. The resulting effective mass becomes very large in high-density regions (suppressing the interaction) while tending to a small (non-zero) value in low-density regions, resulting in an effective fifth force [35] that mimics the influence of a positive cosmological constant in the motion of non-relativistic objects. The scalar field is usually characterized by two parameters: the coupling constant between the scalar field and matter, $\beta$, and the value of the field at infinity, $\phi_{\infty}$, where the matter density matches the background density – the maximum value of the Chameleon field – under reasonable assumptions (e.g. [39]), these parameters describe entirely the modification of gravity. Due to a large amount of high-accuracy observational data, Chameleon theory parameters are tightly constrained, at the laboratory (e.g. [40, 41]), astrophysical (e.g. [42, 43, 44]) and cosmological scales (e.g. [45, 46, 47, 48]), see also [27] for a review. Although the region of the parameter space for viable Chameleon theories has been tightly constrained, there is yet a significant leeway in galaxy cluster scales. The majority of the Chameleon field constraints with clusters were obtained by assuming only the Navarro-Frenk-White (NFW) profile [49] to model the entirety of the cluster’s mass distribution (baryonic+dark matter).
@@
The paper is structured as follows: section 2 reviews the basic theory of the Chameleon gravity and the screening mechanism, further presenting the mass models adopted and the semi-analytical approximation in section 2.2 and section 2.3, respectively. In section 2.2, the solutions for the field profile and its derivative are displayed and tested against the numerical solutions in section 4. section 5 briefly describes the simulation of synthetic clusters and the MG-MAMPOSSt code [52] utilized for the kinematic mass profile reconstruction; the results of the analysis are then presented and discussed. Finally, section 6 summarizes the main conclusions and the key findings.
@@ 2.1 Lagrangian and the equation of motion
with $M_{\rm Pl}=1/\sqrt{8\pi G}$ the reduced Planck mass, while $g_{\mu\nu}^{(i)}=e^{-\frac{2\beta_{i}\phi}{c^{2}M_{\rm Pl}}}\tilde{g}_{\mu\nu}^{(i)}$, with $g_{\mu\nu}^{(i)}$ the metric in the Einstein frame and $\beta_{i}$ the coupling strength of the Chameleon field to each matter field, $\psi^{(i)}$, and $c$ is the speed of light. Following the standard procedure, for simplicity one considers a single effective matter field, $\psi$, that describes the entire matter distribution [^2] While it is possible to extend the analysis to distinct field couplings to the chameleon field, it falls away from the objective of the manuscript. It will be explored in future investigations.. This results in a single coupling constant, $\beta$ [^3] A particular choice for the coupling, $\beta=1/\sqrt{6}$, corresponds to the popular class of $f(\mathcal{R})$ gravity models [53, 54]., and Einstein frame metric $g_{\mu\nu}=e^{-\frac{2\beta\phi}{c^{2}M_{\rm Pl}}}\tilde{g}_{\mu\nu}$. Non-relativistic matter have an energy density $\tilde{\rho}=\rho\,e^{\frac{\beta\phi}{c^{2}M_{\rm Pl}}}$. The potential $V(\phi)$ is a monotonic function of the scalar field.
@@
Solving eq. 2.4 under appropriate boundary conditions for a given mass distribution yields the profile of the Chameleon field, which in turn, originates the fifth force.
In order to describe the mass distribution of a galaxy cluster, one has to resort to a mass density profile model, $\rho$. The standard mass density profile is the NFW model. However, not all galaxy clusters are best described by the latter [^4] Bayesian analysis done in [51], suggests that the three clusters in the GR case prefer either Isothermal and Burket profiles, albeit a weak preference, with the Bayesian evidence $\lesssim 2.5$.. It is then tantamount to extend the analysis to alternative mass density models and generalise the Chameleon screening phenomenology to assess its constraints and thus investigate whether this could be validated in the context of Chameleon gravity or by any means change the evidence.
@@ 2.2 Mass models
$\rho(r)=\frac{\rho_{\rm s}}{\frac{r}{r_{\rm s}}\big{(}1+\frac{r}{r_{\rm s}}\big{)}^{2}}\ ,$ (2.5)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\big{(}\frac{r}{r_{\rm s}}\big{)}\big{(}1+\frac{r}{r_{\rm s}}\big{)}^{b}}\ .$ (2.6)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\big{(}\frac{r}{r_{\rm s}}\big{)}^{\gamma}\big{(}1+\frac{r}{r_{\rm s}}\big{)}^{3-\gamma}}\ .$ (2.7)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\big{(}1+\frac{r}{r_{\rm s}}\big{)}\Big{(}1+\big{(}\frac{r}{r_{\rm s}}\big{)}^{2}\Big{)}}\ .$ (2.8)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\Big{(}\big{(}\frac{r}{r_{\rm s}}\big{)}^{2}+1\Big{)}^{3/2}}\ .$ (2.9)
@@
where $m\in\mathbb{N}$ is a characteristic exponent. While there are additional assumptions that one could utilise for the mass density profiles of galaxy clusters, we find the above selection to be sufficiently extensive for the current analysis: test the effect of the Chameleon mechanism on the shape of the mass distribution.
@@ 2.3 Field solutions
where $C_{\rm s}$ is an integration constant. We now distinguish two cases: $a)$ when the field in the interior region is screened (the aim of this work) and $b)$ when it does not reach the minimum of the effective potential (i.e. no interior solution). Case $b)$ is equivalent to a zero screening radius – or negative –; thus, $\phi(r)$ will not be screened, and the ’exterior’ solution is valid everywhere. The constant of integration in this scenario can be determined by fixing the boundary conditions.
@@ 3.1 Solutions of the NFW-type
Amongst the above equations, the solution to the former, which we term as screening function ( $\rm{f_{s}(r)}$ ) provides us with the screening radius ($r_{\rm c}$). In the case where the field is not screened, the exterior solution holds at any $r>0$, with the integration constant given by,
@@
The unscreened solution is obtained by imposing $C_{\rm s}=-\mathcal{B}$. It is worth pointing out that in eqs. 3.10 and 3.3, $r_{\rm c}$ can be explicitly expressed as a function of $\phi_{\infty}$. Such a relation is, however, not straightforward for all models. Finally, the field profile associated with the $g$NFW matter density comes as:
$\begin{split}\frac{\text{d}\phi_{\text{ext}}}{\text{d}r}=&\frac{\mathcal{B}}{r_{\rm s}x^{2}}\bigg{[}\frac{x^{3-\gamma}}{3-\gamma}\,{{}_{2}}F_{1}(3-\gamma,3-\gamma,4-\gamma,-x)\\ &+1\bigg{]}+\frac{C_{\rm s}}{r_{\rm s}^{2}x^{2}}\ ,\end{split}$ (3.8)
@@
$\begin{split}\phi_{\text{ext}}(x)=&-\frac{\mathcal{B}}{x}\left[\frac{x^{3-\gamma}(1+x)^{\gamma-2}-2-x+\gamma}{\gamma-2}\right.\\ &+\left.\frac{x^{3-\gamma}}{3-\gamma}\,{{}_{2}}F_{1}(3-\gamma,3-\gamma,4-\gamma,-x)\right]\\ &-\frac{C_{\text{s}}}{r_{\text{s}}\,x}+\phi_{\infty}\ .\end{split}$ (3.9)
@@
When the field is in the unscreened regime, no real positive solutions for the screening radius $r_{\rm c}$ can be found. The integration constant of the exterior field eq. 2.14 is then given by $C_{\rm s}=-\mathcal{B}\,r_{\rm s}$.
@@ 3.2 Burkert solutions
$\begin{split}\frac{\text{d}\phi_{\text{ext}}}{\text{d}r}=&\frac{C_{\text{s}}}{r_{\text{s}}\,x^{2}}+\frac{\mathcal{B}}{r_{\text{s}}\,x^{2}}\Big{[}\frac{1}{4}\ln\left(x^{2}+1\right)\\ &+\frac{1}{2}\ln(x+1)-\frac{1}{2}\tan^{-1}(x)\Big{]}\ ,\end{split}$ (3.11)
@@
$C_{\text{s}}=\frac{1}{4}\Big{[}-2\mathcal{B}\log\left(x_{\rm c}^{2}+1\right)+\pi\mathcal{B}-4\phi_{\infty}\Big{]}\ .$
@@
where the matching conditions to get $r_{c}$ cannot be solved analytically. In the left panel of fig. 1 we show the screening function ($\rm{f_{s}(x)}$) plotted for varying values of $\frac{\phi_{\infty}}{\mathcal{B}}$. Wherein each profile has been normalised to its value at $r=0$. The intersection of the screening function profiles at $\rm{f_{s}(r_{\rm c})}=0$ provides the screening radius $r_{\rm c}$ for a given value of $\frac{\phi_{\infty}}{\mathcal{B}}$. It is illustrative to notice that for $\frac{\phi_{\infty}}{\mathcal{B}}\gtrsim 0.785\sim\pi/4$, there exists no solution to the screening function and the entire cluster is unscreened. As it can also be seen by the structure of eq. 3.13, the solution for the screening radius $r_{c}$ demands $\frac{\phi_{\infty}}{\mathcal{B}}\leqslant\pi/4$, i.e. the rhs should be greater or equal to zero since the lhs is a monotonically-increasing function of $x_{c}=r_{c}/r_{s}$, which is zero for $r_{c}=0$, thus, strictly positive for $x_{c}>0$. As mentioned earlier, this corresponds to the case where the field is unscreened and the entire cluster experiences the fifth force.
Figure 1: Variation of the screening function ($\rm{f_{s}(x)}$) as a function of $\frac{\phi_{\infty}}{\mathcal{B}}$, for Burkert (top) and Isothermal (bottom) profiles. The vertical dashed line marks the solution when the screening radius $r_{\rm c}=3r_{\rm s}$ for a given value of $\frac{\phi_{\infty}}{\mathcal{B}}$ for the respective mass models. The intersection of each profile with the $\rm{f_{s}(x)}=0$, provides the value of $r_{\rm c}$, within which the effects of the Chameleon field are screened.
@@ 3.3 Isothermal solutions
Note that in the unscreened regime (i.e. $x_{\rm c}=0$) the constant vanishes, $C_{\text{s}}=0$. In the right panel of fig. 1 we show the screening function profiles for the Isothermal mass density profile corresponding to eq. 3.17. In contrast to the Burkert profile, we find that there always exists a solution to the screening function even in the limit $\frac{\phi_{\infty}}{\mathcal{B}}\to 1$.

### v2

@@ Abstract
The chameleon screening mechanism has been constrained many a time using dynamic and kinematic galaxy cluster observables. Current constraints are, however, insensitive to different mass components within galaxy clusters and have been mainly focused on a single mass density profile, the Navarro-Frenk-White mass density model. In this work, we extend the study of the Chameleon screening mechanism in galaxy clusters by considering a series of mass density models, namely: generalized-Navarro-Frenk-While, b-Navarro-Frenk-While, Burket, Isothermal and Einasto. The coupling strength ($\beta$) and asymptotic value of the chameleon field ($\phi_{\infty}$) are constrained by using kinematics analyses of simulated galaxy clusters, generated both assuming General Relativity and a strong chameleon scenario. By implementing a Bayesian analysis we comprehensively show that the biases introduced due to an incorrect assumption of the mass model are minimal. Similarly, we also demonstrate that a spurious detection of evidence for modifications to gravity is highly unlikely when utilizing the kinematics of galaxy clusters.
@@ 1 Introduction
Among the plethora of possibilities, Chameleon field theories [35, 36, 37] deliver some of the most promising results. In the latter, the environment-dependent effects of gravity are achieved by a non-minimally coupling between an additional scalar field (aka chameleon field) and matter, with the screening achieved by the scalar fields’ potential dependency on the local energy density[38]. The resulting effective mass becomes very large in high-density regions (suppressing the interaction) while tending to a small (non-zero) value in low-density regions, resulting in an effective fifth force [35] that mimics the influence of a positive cosmological constant in the motion of non-relativistic objects. The scalar field is usually characterized by two parameters: the coupling constant between the scalar field and matter, $\beta$, and the value of the field at infinity, $\phi_{\infty}$, where the matter density matches the background density – the maximum value of the Chameleon field – under reasonable assumptions (e.g. [39]), these parameters describe entirely the modification of gravity. Due to a large amount of high-accuracy observational data, Chameleon theory parameters are tightly constrained, at the laboratory (e.g. [40, 41]), astrophysical (e.g. [42, 43, 44, 45]) and cosmological scales (e.g. [46, 47, 48, 49]), see also [27] for a review. Although the region of the parameter space for viable Chameleon theories has been tightly constrained, there is yet a significant leeway in galaxy cluster scales. The majority of the Chameleon field constraints with clusters were obtained by assuming only the Navarro-Frenk-White (NFW) profile [50] to model the entirety of the cluster’s mass distribution (baryonic+dark matter).
@@
The paper is structured as follows: section 2 reviews the basic theory of the Chameleon gravity and the screening mechanism, further presenting the mass models adopted and the semi-analytical approximation in section 2.2 and section 2.3, respectively. In section 3, the solutions for the field profile and its derivative are displayed and tested against the numerical solutions in section 4. section 5 briefly describes the simulation of synthetic clusters and the MG-MAMPOSSt code [53] utilized for the kinematic mass profile reconstruction; the results of the analysis are then presented and discussed. Finally, section 6 summarizes the main conclusions and the key findings.
@@ 2.1 Lagrangian and the equation of motion
with $M_{\rm Pl}=1/\sqrt{8\pi G}$ the reduced Planck mass, while $g_{\mu\nu}^{(i)}=e^{-\frac{2\beta_{i}\phi}{c^{2}M_{\rm Pl}}}\tilde{g}_{\mu\nu}$, with $g_{\mu\nu}^{(i)}$ the metric in the Einstein frame and $\tilde{g}_{\mu\nu}$ the metric in the Jordan frame; $\beta_{i}$ the coupling strength of the Chameleon field to each matter field, $\psi^{(i)}$, and $c$ is the speed of light. Following the standard procedure, for simplicity one considers a single effective matter field, $\psi$, that describes the entire matter distribution [^2] While it is possible to extend the analysis to distinct field couplings to the chameleon field, it falls away from the objective of the manuscript. It will be explored in future investigations.. This results in a single coupling constant, $\beta$ [^3] A particular choice for the coupling, $\beta=1/\sqrt{6}$, corresponds to the popular class of $f(\mathcal{R})$ gravity models [54, 55]., and Einstein frame metric $g_{\mu\nu}=e^{-\frac{2\beta\phi}{c^{2}M_{\rm Pl}}}\tilde{g}_{\mu\nu}$.
Non-relativistic matter have an energy density $\tilde{\rho}=\rho\,e^{\frac{\beta\phi}{c^{2}M_{\rm Pl}}}$. The potential $V(\phi)$ is a monotonic function of the scalar field.
@@
Solving Eq. 2.4 under appropriate boundary conditions – namely, in the absence of a matter-density distribution, the scalar field tends to a constant, background, value, $\phi(r\to\infty)=\phi_{B}$; while at the centre of the system, the matter-density distribution is constant and the scalar field remains suppressed, $\nabla^{2}\phi\approx 0$ – for a given mass distribution yields the profile of the Chameleon field, which in turn, originates the fifth force.
In order to describe the mass distribution of a galaxy cluster, one has to resort to a mass density profile, $\rho$. So far, the majority of the studies of Chameleon gravity at cluster scales in the literature relied on the standard NFW density model. However, not all galaxy clusters are best described by the latter [^4] Bayesian analysis done in [52], suggests that the three clusters in the GR case prefer either Isothermal and Burket profiles, albeit a weak preference, with the change in Bayesian evidence $\lesssim 2.5$.. It is then tantamount to extend the analysis to alternative mass density models, and explore the respective Chameleon screening phenomenology in order to assess its constraints and thus investigate whether such mass models could be validated in the context of Chameleon gravity or by any means change the evidence.
@@ 2.2 Mass models
$\rho(r)=\frac{\rho_{\rm s}}{\frac{r}{r_{\rm s}}\left(1+\frac{r}{r_{\rm s}}\right)^{2}}\ ,$ (2.5)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\left(\frac{r}{r_{\rm s}}\right)\left(1+\frac{r}{r_{\rm s}}\right)^{b}}\ .$ (2.6)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\left(\frac{r}{r_{\rm s}}\right)^{\gamma}\left(1+\frac{r}{r_{\rm s}}\right)^{3-\gamma}}\ .$ (2.7)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\left(1+\frac{r}{r_{\rm s}}\right)\left[1+\Big{(}\frac{r}{r_{\rm s}}\Big{)}^{2}\right]}\ .$ (2.8)
@@
$\rho(r)=\frac{\rho_{\rm s}}{\Big{[}\left(\frac{r}{r_{\rm s}}\right)^{2}+1\Big{]}^{3/2}}\ .$ (2.9)
@@
where $m\in\mathbb{N}$ is a characteristic exponent.
In Fig. 1 the density profiles for all the models introduced above are shown for a reference value of the critical density $\rho_{\rm s}=5\times 10^{14}M_{\odot}/\text{Mpc}^{3}$ and a scale radius $r_{\rm s}=0.5\,\text{Mpc}$. The Isothermal, Burkert and Einasto profiles flatten to a constant value for $r\to 0$, while the NFW and its generalizations exhibit a cusp, diverging at small radii. Increasing $b$ in the b-NFW models provides a faster suppression of the density at large $r$, whereas an increase of the $m$ parameter in the Einasto model results in a shallower profile. In total, the set of density models captures a quite broad range of behaviours.
Figure 1: Density profile models used in this work, for a halo with $\rho_{\rm s}=5\times 10^{14}M_{\odot}/\text{Mpc}^{3}$ and $r_{\rm s}=0.5\,\text{Mpc}$. Left: Burkert, NFW, Hernquist (b-NFW with $b=3$); Right: gNFW with $\gamma=1.2$, Isothermal and Einasto with $m=2$.
While there are additional assumptions that one could utilise for the mass density profiles of galaxy clusters, we find the above selection to be sufficiently extensive for the current analysis: test the effect of the Chameleon mechanism on the shape of the mass distribution.
@@ 2.3 Field solutions
where $C_{\rm s}$ is an integration constant. We now distinguish two cases: $a)$ when the field in the interior region is screened (the aim of this work) and $b)$ when it does not reach the minimum of the effective potential (i.e. no interior solution). Case $b)$ is equivalent to a zero screening radius – or negative – thus, $\phi(r)$ will not be screened, and the ’exterior’ solution is valid everywhere. The constant of integration in this scenario can be determined by fixing the boundary conditions.
@@ 3.1 Solutions of the NFW-type
Amongst the above equations, the solution to the former, which we term as screening function ($\rm{f_{s}(r)}$), provides us with the screening radius ($r_{\rm c}$). In the case where the field is not screened, the exterior solution holds at any $r>0$, with the integration constant given by,
@@
The unscreened solution is obtained by imposing $C_{\rm s}=-\mathcal{B}$. Finally, the field profile associated with the $g$NFW matter density comes as:
$\begin{split}\frac{\text{d}\phi_{\text{ext}}}{\text{d}r}=\frac{\mathcal{B}}{r_{\rm s}x^{2}}\bigg{[}\frac{x^{3-\gamma}}{3-\gamma}\,{{}_{2}}F_{1}(3-\gamma,3-\gamma,4-\gamma,-x)+1\bigg{]}+\frac{C_{\rm s}}{r_{\rm s}^{2}x^{2}}\ ,\end{split}$ (3.8)
@@
$\begin{split}\phi_{\text{ext}}(x)=-\frac{\mathcal{B}}{x}\left[\frac{x^{3-\gamma}(1+x)^{\gamma-2}-2-x+\gamma}{\gamma-2}+\frac{x^{3-\gamma}}{3-\gamma}\,{{}_{2}}F_{1}(3-\gamma,3-\gamma,4-\gamma,-x)\right]-\frac{C_{\text{s}}}{r_{\text{s}}\,x}+\phi_{\infty}\ .\end{split}$ (3.9)
@@
When the field is in the unscreened regime, no real positive solutions for the screening radius $r_{\rm c}$ can be found. The integration constant of the exterior field Eq. 2.14 is then given by $C_{\rm s}=-\mathcal{B}\,r_{\rm s}$. It is worth pointing out that in Eqs. 3.10 and 3.3, $r_{\rm c}$ can be explicitly expressed as a function of $\phi_{\infty}$. Such a relation is, however, not straightforward for all models.
@@ 3.2 Burkert solutions
$\begin{split}\frac{\text{d}\phi_{\text{ext}}}{\text{d}r}=\frac{C_{\text{s}}}{r_{\text{s}}\,x^{2}}+\frac{\mathcal{B}}{r_{\text{s}}\,x^{2}}\Big{[}\frac{1}{4}\ln\left(x^{2}+1\right)+\frac{1}{2}\ln(x+1)-\frac{1}{2}\tan^{-1}(x)\Big{]}\ ,\end{split}$ (3.11)
@@
$C_{\text{s}}=\frac{1}{4}\mathcal{B}\left[-\log\left(x_{\rm c}^{2}+1\right)-2\log(x_{\rm c}+1)+2\tan^{-1}(x_{\rm c})\right]$
@@
where the matching conditions to get $r_{c}$ cannot be solved analytically. In the top panel of Fig. 2 we show the screening function $\rm{f_{s}}(x_{\rm c})/{\rm f_{s}}(0)$, with
${\rm f_{s}}(x_{\rm c})=\ln\left[\frac{x_{\rm c}^{2}+1}{(x_{\rm c}+1)^{2}}\right]+2\tan^{-1}(x_{\rm c})-\pi+4\frac{\phi_{\infty}}{\mathcal{B}}\,,$
plotted for varying values of $\frac{\phi_{\infty}}{\mathcal{B}}$. Wherein each profile has been normalised to its value at $r=0$. The intersection of the screening function profiles at $\rm{f_{s}(r_{\rm c})}=0$ provides the screening radius $r_{\rm c}$ for a given value of $\frac{\phi_{\infty}}{\mathcal{B}}$. It is illustrative to notice that for $\frac{\phi_{\infty}}{\mathcal{B}}\gtrsim 0.785\sim\pi/4$, there exists no solution to the screening function and the entire cluster is unscreened. As it can also be seen by the structure of Eq. 3.13, the solution for the screening radius $r_{c}$ demands $\frac{\phi_{\infty}}{\mathcal{B}}\leqslant\pi/4$, i.e. the rhs should be greater or equal to zero since the lhs is a monotonically-increasing function of $x_{c}=r_{c}/r_{s}$, which is zero for $r_{c}=0$, thus, strictly positive for $x_{c}>0$. As mentioned earlier, this corresponds to the case where the field is unscreened and the entire cluster experiences the fifth force.
Figure 2: Variation of the screening function ($\rm{f_{s}(x)}$) as a function of $\frac{\phi_{\infty}}{\mathcal{B}}$, for Burkert (top) and Isothermal (bottom) profiles. The vertical dashed line marks the solution when the screening radius $r_{\rm c}=3r_{\rm s}$ for a given value of $\frac{\phi_{\infty}}{\mathcal{B}}$ for the respective mass models. The intersection of each profile with the $\rm{f_{s}(x)}=0$, provides the value of $r_{\rm c}$, within which the effects of the Chameleon field are screened. To avoid the sign ambiguity of the screening function we show the re-scaled $\text{f}_{\text{s}}(r)/\text{f}_{\text{s}}(r=0)$.
@@ 3.3 Isothermal solutions
Note that in the unscreened regime (i.e. $x_{\rm c}=0$) the constant vanishes, $C_{\text{s}}=0$. In the bottom panel of Fig. 2 we show the screening function profiles for the Isothermal mass density profile corresponding to Eq. 3.17, $\text{f}_{\text{s}}(x_{\rm c})=\sqrt{x_{\rm c}^{2}+1}-\mathcal{B}/\phi_{\infty}$. In contrast to the Burkert profile, we find that there always exists a solution to the screening function even in the limit $\frac{\phi_{\infty}}{\mathcal{B}}\to 1$.